OpenStudy (anonymous):

find all values of x in the interval[0,2pie] that satisfies the equation sin2x=cosx

OpenStudy (rulnick):

pi/2 and 3pi/2

OpenStudy (ash2326):

We have \[\sin 2x= \cos x\] \[2 \sin x \cos x-\cos x=0\] so \[\cos x(2 \sin x -1)=0\] either cos x=0 or 2 sin x-1=0=> sin x =1/2 cos x=0 \[x= \pi/2, 3\pi/2\] sin x=1/2 \[x=\pi/6, (\pi-\pi /6)=>\pi/6, 5\pi /6\] so \[x= \pi/6, \pi/2, 5\pi/6, 3\pi/2\]

OpenStudy (anonymous):

this is because the interval is [0,2pi] right

OpenStudy (ash2326):

yeah, otherwise there'd be many solutions

OpenStudy (anonymous):

ok again thank you so much :)

OpenStudy (rulnick):

oops, I read it as sin(2x)=cos(x)=0, missed the two other sol'ns. good job getting the other two.